# Stability of $L^{2}-$invariants on stratified spaces

### v2

@@ 1.2. Statement of the main results
Let $\overline{M}$ and $\overline{N}$ be two compact smoothly stratified pseudo-manifolds which satisfy Assumption 3.2; assume that there exist two Galois $\Gamma$-coverings $\overline{M}_{\Gamma}$, $\overline{N}_{\Gamma}$ and a smoothly stratified codimension-preserving homotopy equivalence $f:\overline{M}\to\overline{N}$ satisfying condition (1.2) (see (4.1) for precise notation). Then
$\begin{split}b^{*}_{(2),\min/\max}(\overline{M}_{\Gamma})&=b^{*}_{(2),\min/\max}(\overline{N}_{\Gamma}),\\ \alpha_{*,\min/\max}(\overline{M}_{\Gamma})&=\alpha_{*,\min/\max}(\overline{N}_{\Gamma}).\end{split}$ (1.3)
@@
Let $\overline{M}$ and $\overline{N}$ be two compact smoothly stratified pseudo-manifolds which satisfy Assumption 3.2; assume that there exists a smoothly stratified codimension-preserving homotopy equivalence $f:\overline{M}\to\overline{N}$ which satisfies the condition (4.1). Then
$\begin{split}b^{*}_{(2),\min/\max}(\overline{M})&=b^{*}_{(2),\min/\max}(\overline{N}),\\ \alpha_{*,\min/\max}(\overline{M})&=\alpha_{*,\min/\max}(\overline{N})\end{split}$ (1.4)
where $b^{*}_{(2),\min/\max}(\overline{M})$, $b^{*}_{(2),\min/\max}(\overline{N})$, $\alpha_{*,\min/\max}(\overline{M})$ and $\alpha_{*,\min/\max}(\overline{N})$ denote the maximal/minimal $L^{2}$-Betti numbers and Novikov-Shubin invariants of $\overline{M}$ and $\overline{N}$ computed with respect to the corresponding universal coverings.
@@ 1.3. Background and notation
    A Galois covering $\pi:\overline{M}_{\Gamma}\to\overline{M}$ with Galois group $\Gamma$ is again a smoothly stratified (Thom-Mather) pseudo-manifold with the smooth open and dense stratum $M_{\Gamma}$ being a Galois covering of $M$. Any singular stratum $Y_{\Gamma}$ of $\overline{M}_{\Gamma}$ is a Galois covering of a singular stratum $Y$ in $\overline{M}$ of same depTheorem The lift $g_{\Gamma}$ of the wedge metric $g$ defines a $\Gamma$-invariant wedge metric on $M_{\Gamma}$. We refer to e.g. [PiVe19, §7.1] and [PiVe22, §2.4] for more details.
@@
Acknowledgements. The authors would like to acknowledge interesting discussions with Pierre Albin and Matthias Lesch. The third author thanks Sapienza University for hospitality. This work was supported by the ”National Group for the Algebraic and Geometric Structures and their Applications” (GNSAGA-INDAM).
@@ 2. Reformulating the problem using Mishchenko bundles
In fact, [Sch05, Lemma 7.10] proves that these two spaces are $\mathscr{N}\Gamma$-Hilbert modules in the sense of [Sch05, Definition 2.1], and as explained in [Sch05, §7.5, (2)], the isomorphism $\Phi$ extends to an isometry of $\mathscr{N}\Gamma$-Hilbert modules
@@ 3.1. Some analytic properties
$D_{\textup{rel}}:=d_{\Gamma,\,\min}+d^{*}_{\Gamma,\,\min},\quad\Delta_{\Gamma,\mathrm{rel}}:=D^{2}_{\Gamma,\mathrm{rel}},$
$D_{\textup{abs}}:=d_{\Gamma,\,\max}+d^{*}_{\Gamma,\,\max},\quad\Delta_{\Gamma,\mathrm{abs}}:=D^{2}_{\Gamma,\mathrm{abs}}.$
@@
Let us denote with $\Psi:L^{2}\Omega^{*}(\mathrm{reg}(U),g_{\Gamma}|_{\mathrm{reg}(U)})\rightarrow L^{2}\Omega^{*}(\mathrm{reg}(V),g|_{\mathrm{reg}(V)})$ the isometry induced by $(p|_{U})^{-1}:V\rightarrow U$. Here, we write $\mathrm{reg}(U)$ and $\mathrm{reg}(V)$ for the open interior of $U$ and $V$, respectively. Let $\phi\in\mathcal{A}(\overline{M}_{\Gamma})$ with $\mathrm{supp}(\phi)\subset U$ and $\omega\in\mathscr{D}(D_{\Gamma,\mathrm{abs}/\mathrm{rel}})$ be arbitrarily fixed. It is not difficult to show that
@@ 4.2. Construction
The corresponding pullbacks, defined by ${}^{e}df$ and ${}^{w}df$, consequently act as follows
@@
where $\mathscr{E}$ is the Mishchenko bundle over $\overline{N}$, as in (2.1), pulled back to $\widetilde{N}$ by the blowdown map $\beta:\widetilde{N}\to\overline{N}$ that is associated with the resolution $\widetilde{N}$.
@@
We define the $e$-differential ${}^{e}df$ as follows. By assumption in Definition 4.7, $f$ preserves the iterated boundary fibrations structures and hence sends curves that are tangent to the fibres at the boundary of $M$ to curves that are tangent to the fibres at the boundary of $N$. Hence we obtain ${}^{e}df:{}^{e}TM\to{}^{e}TN$.
For the action of the differential between the wedge tangent bundles, recall that the total boundary functions $\rho_{M}\in C^{\infty}(\widetilde{M})$ and $\rho_{N}\in C^{\infty}(\widetilde{N})$ are nowhere vanishing in the open interior, and vanish to first order at each boundary face of $\widetilde{M}$ and $\widetilde{N}$, respectively. We write for any $V\in T_{p}M$
@@
If $V$ is now a section of ${}^{w}TM$, $\rho_{M}V$ extends to a section of ${}^{e}TM$, and $df(\rho_{M}V)$ extends to a section ${}^{e}df(\rho_{M}V)$ of ${}^{e}TN$. Consequently, $\rho^{-1}_{N}\circ f\cdot df(V)$ extends to a section $\widetilde{f}^{*}\rho^{-1}_{N}\cdot{}^{e}df(\rho_{M}V)$ of ${}^{w}TN$. Now, since $f$ is smoothly stratified, we have that $\widetilde{f}$ sends boundary hypersurfaces of $\widetilde{M}$ onto boundary hypersurfaces of $\widetilde{N}$. Thus $\rho^{-1}_{M}\cdot\widetilde{f}^{*}\rho_{N}$ is smooth and bounded (note that this is not the case for an arbitrary $b$-map), and we can define in view of (4.3) for any $V\in{}^{w}TM$
${}^{w}df(V):=\frac{\widetilde{f}^{*}\rho_{N}}{\rho_{M}}\cdot\widetilde{f}^{*}\rho^{-1}_{N}\cdot{}^{e}df(\rho_{M}V)\in{}^{w}TN.$ (4.4)
The action of the pullbacks ${}^{e}f^{*}$ and ${}^{w}f^{*}$, defined by ${}^{e}df$ and ${}^{w}df$, respectively, is then obvious. ∎
@@
($\begin{split}&{}^{e}(\exp\circ\,\mathbb{B}\widetilde{f})^{*}:C^{\infty}(\widetilde{N},\Lambda^{*}({}^{e}T^{*}\widetilde{N})\otimes\mathscr{E})\to C^{\infty}(\widetilde{M},\Lambda^{*}({}^{e}T^{*}\widetilde{f}^{*}(\mathbb{B}^{N})\otimes\mathscr{E}^{\prime}),\\ &{}^{w}(\exp\circ\,\mathbb{B}\widetilde{f})^{*}:C^{\infty}(\widetilde{N},\Lambda^{*}({}^{w}T^{*}\widetilde{N})\otimes\mathscr{E})\to C^{\infty}(\widetilde{M},\Lambda^{*}({}^{w}T^{*}\widetilde{f}^{*}(\mathbb{B}^{N})\otimes\mathscr{E}^{\prime}).\end{split}$ (4.6)
@@
for $t\in[0,1]$. Pullback by homotopic maps gives isomorphic bundles and hence there exists an isomorphism $\mathscr{J}^{\prime}:r_{0}^{*}\exp^{*}\mathscr{E}\to r_{1}^{*}\exp^{*}\mathscr{E}$. Consider the flat connection $\nabla$ on $\exp^{*}\mathscr{E}$, obtained as the pullback of the canonical flat connection on the Mishchenko bundle $\mathscr{E}$. We claim that this isomorphism $\mathscr{J}^{\prime}$ commutes with the pullback connections $r_{0}^{*}\nabla\equiv\nabla$ and $r_{1}^{*}\nabla$. Indeed, consider
@@
.$*.\displaystyle\begin{array}[]{lll}&\Bigl{.}r^{*}\exp^{*}\mathscr{E}\Bigr{|}_{\{0\}\times\mathbb{B}^{N}}=r_{0}^{*}\exp^{*}\mathscr{E}\equiv\exp^{*}\mathscr{E},&\Bigl{.}r^{*}\nabla\Bigr{|}_{\{0\}\times\mathbb{B}^{N}}=r_{0}^{*}\nabla\equiv\nabla,\\ &\Bigl{.}r^{*}\exp^{*}\mathscr{E}\Bigr{|}_{\{1\}\times\mathbb{B}^{N}}=r_{1}^{*}\exp^{*}\mathscr{E},&\Bigl{.}r^{*}\nabla\Bigr{|}_{\{1\}\times\mathbb{B}^{N}}=r_{1}^{*}\nabla.\end{array}$
The isomorphism $\mathscr{J}^{\prime}$ may be constructed via parallel transport $P$ by $r^{*}\nabla$: given any $(0,p)\in[0,1]\times\mathbb{B}^{N}$, consider the path $\gamma(t):=(t,p),t\in[0,1]$ and define
@@
It is an isometry, since the metric in each fibre is given by the $\ell^{2}\Gamma$ inner product and parallel transport preserves the inner product. The isomorphism commutes with the pullback connections $r_{0}^{*}\nabla$ and $r_{1}^{*}\nabla$, since it maps parallel sections to parallel sections: indeed, for any path $\gamma^{\prime}$ between $p,q,\in\mathbb{B}^{N}$ we have as indicated in Figure 5
$P_{\gamma}\circ P_{\gamma^{\prime}}=P_{\gamma^{\prime}}\circ P_{\gamma},$
{tikzpicture}
Figure 5. Parallel transport across $[0,1]\times\mathbb{B}^{N}$.
since parallel transport for flat connections is trivial along closed paths. Thus, parallel sections along $\gamma^{\prime}$ with respect to $r_{0}^{*}\nabla$ are mapped by $\mathscr{J}^{\prime}$ to parallel sections along $\gamma^{\prime}$ with respect to $r_{1}^{*}\nabla$. This proves our claim and we can now finish the proof.
We have the following sequence of bundle isomorphisms (note that $r_{1}=\textup{id}$)
$\exp^{*}\mathscr{E}=r_{1}^{*}\exp^{*}\mathscr{E}\cong r_{0}^{*}\exp^{*}\mathscr{E}=(\exp\,\circ\,r_{0})^{*}\mathscr{E}=(\pi_{N}\circ r_{0})^{*}\mathscr{E}=\pi_{N}^{*}\mathscr{E}.$
@@
We can now define the Hilsum-Skandalis replacement for $f^{*}$.

### v3

@@ 1.2. Statement of the main results
Let $\overline{M}$ and $\overline{N}$ be oriented compact smoothly stratified pseudo-manifolds which satisfy Assumption 3.2; assume that there exist two Galois $\Gamma$-coverings $\overline{M}_{\Gamma}$, $\overline{N}_{\Gamma}$ and a smoothly stratified codimension-preserving homotopy equivalence $f:\overline{M}\to\overline{N}$ satisfying condition (1.2) (see (4.1) for precise notation). Then
$b^{*}_{(2),\max}(\overline{M}_{\Gamma})$ $=b^{*}_{(2),\max}(\overline{N}_{\Gamma}),\quad b^{*}_{(2),\min}(\overline{M}_{\Gamma})=b^{*}_{(2),\min}(\overline{N}_{\Gamma}),$
$\alpha_{*,\max}(\overline{M}_{\Gamma})$ $=\alpha_{*,\max}(\overline{N}_{\Gamma}),\qquad\alpha_{*,\min}(\overline{M}_{\Gamma})=\alpha_{*,\min}(\overline{N}_{\Gamma}).$
As a consequence of Corollary 4.4, the condition (4.1) is always true when we deal with universal coverings. Therefore we have the following result.
@@
Let $\overline{M}$ and $\overline{N}$ be two oriented compact smoothly stratified pseudo-manifolds which satisfy Assumption 3.2; assume that there exists a smoothly stratified codimension-preserving homotopy equivalence $f:\overline{M}\to\overline{N}$. Let $b^{*}_{(2),\min/\max}(\overline{M})$, $b^{*}_{(2),\min/\max}(\overline{N})$, $\alpha_{*,\min/\max}(\overline{M})$ and $\alpha_{*,\min/\max}(\overline{N})$ denote the maximal/minimal $L^{2}$-Betti numbers and Novikov-Shubin invariants of $\overline{M}$ and $\overline{N}$ computed with respect to the corresponding universal coverings. We then have
$b^{*}_{(2),\max}(\overline{M})$ $=b^{*}_{(2),\max}(\overline{N}),\quad b^{*}_{(2),\min}(\overline{M})=b^{*}_{(2),\min}(\overline{N}),$
$\alpha_{*,\max}(\overline{M})$ $=\alpha_{*,\max}(\overline{N}),\qquad\alpha_{*,\min}(\overline{M})=\alpha_{*,\min}(\overline{N}).$
@@
Remark 1.4.
For smooth compact manifolds it is possible to give a topological definition of the $L^{2}$-Betti numbers and of the Novikov-Shubin invariants. The proof of the equivalence of these two definitions is a non-trivial result, treated in great detail in [Lue02]. We believe that it should be possible to employ $L^{2}$-intersection chains à la Goresky-MacPherson on the universal covering of a Witt space in order to give a topological definition of the invariants treated in this article. While the definition does not pose particular problems, the proof of the equivalence between the analytic definition given here and this topological definition appears to be rather involved. We plan to look into these questions in a subsequent paper.
@@ 1.3. Background and notation
    A Galois covering $\pi:\overline{M}_{\Gamma}\to\overline{M}$ with Galois group $\Gamma$ is again a smoothly stratified (Thom-Mather) pseudo-manifold with the smooth open and dense stratum $M_{\Gamma}$ being a Galois covering of $M$. Any singular stratum $Y_{\Gamma}$ of $\overline{M}_{\Gamma}$ is a Galois covering of a singular stratum $Y$ in $\overline{M}$ of same depth. The lift $g_{\Gamma}$ of the wedge metric $g$ defines a $\Gamma$-invariant wedge metric on $M_{\Gamma}$. We refer to e.g. [PiVe19, §7.1] and [PiVe22, §2.4] for more details.
@@
Acknowledgements. The authors would like to acknowledge interesting discussions with Pierre Albin and Matthias Lesch. The third author thanks Sapienza University for hospitality. This work was supported by the ”National Group for the Algebraic and Geometric Structures and their Applications” (GNSAGA-INDAM). We thank the referee for useful comments and questions that certainly improved this article.
@@ 2. Reformulating the problem using Mishchenko bundles
In fact, [Sch05, Lemma 7.10] proves that these two spaces are $\mathscr{N}\Gamma$-Hilbert modules in the sense of [Sch05, Definition 2.1], and as explained in [Sch05, §7.5, (2)], the isomorphism $\Phi_{c}$ extends to an isometry of $\mathscr{N}\Gamma$-Hilbert modules
@@ 3.1. Some analytic properties
$D_{\Gamma,\textup{rel}}:=d_{\Gamma,\,\min}+d^{*}_{\Gamma,\,\min},\quad\Delta_{\Gamma,\mathrm{rel}}:=D^{2}_{\Gamma,\mathrm{rel}},$
$D_{\Gamma,\textup{abs}}:=d_{\Gamma,\,\max}+d^{*}_{\Gamma,\,\max},\quad\Delta_{\Gamma,\mathrm{abs}}:=D^{2}_{\Gamma,\mathrm{abs}}.$
@@
Let us denote with $\Psi:L^{2}\Omega^{*}(\mathrm{reg}(U),g_{\Gamma}|_{\mathrm{reg}(U)})\rightarrow L^{2}\Omega^{*}(\mathrm{reg}(V),g|_{\mathrm{reg}(V)})$ the isometry induced by $(p|_{U})^{-1}:V\rightarrow U$. Here, we write $\mathrm{reg}(U)$ and $\mathrm{reg}(V)$ for the regular part of $U$ and $V$, respectively. Let $\phi\in\mathcal{A}(\overline{M}_{\Gamma})$ with $\mathrm{supp}(\phi)\subset U$ and $\omega\in\mathscr{D}(D_{\Gamma,\mathrm{abs}/\mathrm{rel}})$ be arbitrarily fixed. It is not difficult to show that
@@ 4.2. Construction
The corresponding pullbacks are defined by ${}^{e}df$ and ${}^{w}df$ pointwise as follows.
  1. (1)
    for any $V_{1},\cdots V_{k}\in{}^{e}TM$ and $\omega\in C^{\infty}(\widetilde{N},\Lambda^{*}({}^{e}T^{*}N)\otimes\mathscr{E})$,
    ${}^{e}f^{*}\omega(V_{1},\cdots V_{k}):=\omega({}^{e}df[V_{1}],\cdots{}^{e}df[V_{k}])\circ f.$ (4.2)
  2. (2)
    for any $V_{1},\cdots V_{k}\in{}^{w}TM$ and $\omega\in C^{\infty}(\widetilde{N},\Lambda^{*}({}^{w}T^{*}N)\otimes\mathscr{E})$,
    ${}^{w}f^{*}\omega(V_{1},\cdots V_{k}):=\omega({}^{w}df[V_{1}],\cdots{}^{w}df[V_{k}])\circ f.$ (4.3)
The pullbacks consequently act as follows
@@
where $\mathscr{E}$ is the Mishchenko bundle over $\overline{N}$, as in (2.1), pulled back to $\widetilde{N}$ by the blowdown map $\beta:\widetilde{N}\to\overline{N}$ that is associated with the resolution $\widetilde{N}$. Moreover, the pullbacks commute with the twisted differentials $d_{\mathscr{E}}$ and $d_{\widetilde{f}^{*}\mathscr{E}}$ on the respective spaces of differential forms.
@@
The continuous extension of the differential to the edge tangent bundle, the so-called $e$-differential ${}^{e}df$, is obtained as follows. By assumption in Definition 4.7, $f$ preserves the iterated boundary fibrations structures and hence sends curves that are tangent to the fibres at the boundary of $M$ to curves that are tangent to the fibres at the boundary of $N$. Hence we obtain ${}^{e}df:{}^{e}TM\to{}^{e}TN$.
For the action of the differential between the wedge tangent bundles, recall that the total boundary functions $\rho_{M}\in C^{\infty}(\widetilde{M})$ and $\rho_{N}\in C^{\infty}(\widetilde{N})$ are nowhere vanishing in the open interior, and vanish to first order at each boundary face of $\widetilde{M}$ and $\widetilde{N}$, respectively. We write for any $V\in T_{p}M$ with $p\in M$ in the open interior
@@
If $V$ is now a section of ${}^{w}TM$, $\rho_{M}V$ extends to a section of ${}^{e}TM$, and $df(\rho_{M}V)$ extends to a section ${}^{e}df(\rho_{M}V)$ of ${}^{e}TN$. Consequently, $(\rho^{-1}_{N}\circ\widetilde{f})\cdot df(V)$ extends to a section $\widetilde{f}^{*}\rho^{-1}_{N}\cdot{}^{e}df(\rho_{M}V)$ of ${}^{w}TN$. Now, since $f$ is smoothly stratified, we have that $\widetilde{f}$ sends boundary hypersurfaces of $\widetilde{M}$ onto boundary hypersurfaces of $\widetilde{N}$. Thus, as already explained in [ALMP12, p. 298], $\rho^{-1}_{M}\cdot\widetilde{f}^{*}\rho_{N}$ is smooth and bounded (note that this is not the case for an arbitrary $b$-map), and we can write in view of (4.5) for any $V\in{}^{w}TM$
${}^{w}df(V)=\frac{\widetilde{f}^{*}\rho_{N}}{\rho_{M}}\cdot\widetilde{f}^{*}\rho^{-1}_{N}\cdot{}^{e}df(\rho_{M}V)\in{}^{w}TN.$ (4.6)
This provides a continuous extension of the differential to the wedge tangent bundles. If fact, all the boundary defining functions in (4.6) combine together to the constant function $1$, but the peculiar expression is needed to explain that we indeed have a continuous extension to wedge vector fields in ${}^{w}TM$.
The action of the pullbacks ${}^{e}f^{*}$ and ${}^{w}f^{*}$, as defined by ${}^{e}df$ and ${}^{w}df$ in (4.2) and (4.3) respectively, now follows. It remains to prove that the pullbacks commute with the differentials. Let us write out the argument for the wedge case, the edge case is treated exactly in the same manner. Consider any $V_{0},\ldots,V_{k}\in{}^{w}TM$, $U_{0},\ldots,U_{k}\in{}^{w}TN$ and $\omega\in C^{\infty}(\widetilde{N},\Lambda^{k}({}^{w}T^{*}N)\otimes\mathscr{E})$. Locally, $\omega$ can be written as a sum of differential forms in $C^{\infty}(\widetilde{N},\Lambda^{k}({}^{w}T^{*}N))$ tensorized by parallel sections of $\mathscr{E}$. For such terms we compute as follows.
$d_{\mathscr{E}}\omega(U_{0},\ldots U_{k})$ $=\sum_{i=0}^{\infty}(-1)^{i}U_{i}\Bigl{\{}\omega\left(U_{0},\ldots,\widehat{U_{i}},\ldots,U_{k}\right)\Bigr{\}},$
${}^{w}f^{*}d_{\mathscr{E}}\omega\left(V_{0},\ldots,V_{k}\right)$ $=d_{\mathscr{E}}\omega\bigl{(}{}^{w}df[V_{0}],\ldots,{}^{w}df[V_{k}]\bigr{)}\circ f$
$=\sum_{i=0}^{\infty}(-1)^{i}\ {}^{w}df[V_{i}]\Bigl{\{}\omega\left({}^{w}df[V_{0}],\ldots,\widehat{{}^{w}df[V_{i}]},\ldots,{}^{w}df[V_{k}]\right)\Bigr{\}}\circ f$
$=\sum_{i=0}^{\infty}(-1)^{i}V_{i}\Bigl{\{}\omega\left({}^{w}df[V_{0}],\ldots,\widehat{{}^{w}df[V_{i}]},\ldots,{}^{w}df[V_{k}]\right)\circ f\Bigr{\}}$
$=d_{\widetilde{f}^{*}\mathscr{E}}{}^{w}f^{*}\omega\bigl{(}V_{0},\ldots,V_{k}\bigr{)}.$
@@
($\begin{split}&{}^{e}(\exp\circ\,\mathbb{B}\widetilde{f})^{*}:C^{\infty}(\widetilde{N},\Lambda^{*}({}^{e}T^{*}\widetilde{N})\otimes\mathscr{E})\to C^{\infty}(\widetilde{f}^{*}(\mathbb{B}^{N}),\Lambda^{*}({}^{e}T^{*}\widetilde{f}^{*}(\mathbb{B}^{N})\otimes\mathscr{E}^{\prime}),\\ &{}^{w}(\exp\circ\,\mathbb{B}\widetilde{f})^{*}:C^{\infty}(\widetilde{N},\Lambda^{*}({}^{w}T^{*}\widetilde{N})\otimes\mathscr{E})\to C^{\infty}(\widetilde{f}^{*}(\mathbb{B}^{N}),\Lambda^{*}({}^{w}T^{*}\widetilde{f}^{*}(\mathbb{B}^{N})\otimes\mathscr{E}^{\prime}).\end{split}$ (4.8)
@@
for $t\in[0,1]$. Clearly pullback by homotopic maps gives isomorphic bundles. We use now the homotopy and the flat connection to explicitly construct a bundle isomorphism below, denoted by $\mathscr{J}^{\prime}$, that preserves the given flat connections
$\mathscr{J}^{\prime}:r_{0}^{*}\exp^{*}\mathscr{E}\to r_{1}^{*}\exp^{*}\mathscr{E}.$
Consider the flat connection $\nabla$ on $\exp^{*}\mathscr{E}$, obtained as the pullback of the canonical flat connection on the Mishchenko bundle $\mathscr{E}$. We claim that this isomorphism $\mathscr{J}^{\prime}$ commutes with the pullback connections $r_{0}^{*}\nabla$ and $r_{1}^{*}\nabla$. Indeed, consider
@@
.$\begin{array}[]{lll}&\Bigl{.}r^{*}\exp^{*}\mathscr{E}\Bigr{|}_{\{0\}\times\mathbb{B}^{N}}=r_{0}^{*}\exp^{*}\mathscr{E},&\Bigl{.}r^{*}\nabla\Bigr{|}_{\{0\}\times\mathbb{B}^{N}}=r_{0}^{*}\nabla,\\ &\Bigl{.}r^{*}\exp^{*}\mathscr{E}\Bigr{|}_{\{1\}\times\mathbb{B}^{N}}=r_{1}^{*}\exp^{*}\mathscr{E},&\Bigl{.}r^{*}\nabla\Bigr{|}_{\{1\}\times\mathbb{B}^{N}}=r_{1}^{*}\nabla.\end{array}$
The isomorphism $\mathscr{J}^{\prime}$ may be constructed explicitly via parallel transport $P$ by $r^{*}\nabla$: given any $(0,p)\in[0,1]\times\mathbb{B}^{N}$, consider the path $\gamma(t):=(t,p),t\in[0,1]$ and define
@@
It is an isometry, since the metric in each fibre is given by the $\ell^{2}\Gamma$ inner product and parallel transport preserves the inner product. The isomorphism commutes with the pullback connections $r_{0}^{*}\nabla$ and $r_{1}^{*}\nabla$, if for any continuous path $\gamma^{\prime}$ between $p,q,\in\mathbb{B}^{N}$
$P_{\gamma}\circ P_{\gamma^{\prime}}=P_{\gamma^{\prime}}\circ P_{\gamma}.$ (4.9)
This is illustrated in the commutative diagram in Figure 5. We illustrate the curves $\gamma$ and $\gamma^{\prime}$ as well as the parallel transports along those in Figure 6.
[…]
Figure 5. Parallel transport commutes along null-homotopic paths.
Figure 6. Illustration of parallel transports across $[0,1]\times\mathbb{B}^{N}$.
Now (4.9) holds since parallel transport for flat connections is trivial along null-homotopic paths. Let us explain why the closed path $F$ in Figure 6 is null-homotopic. The path $F$ is by definition traversed by travelling along the following sequence of paths in $[0,1]\times\mathbb{B}^{N}$
  1. (1)
    $(0,\gamma^{\prime}(t))\in\{0\}\times\mathbb{B}^{N},\ t\in[0,1]$
    (we move along the bottom red curve at the left hand side in Fig. 6),
  2. (2)
    $(t,q=\gamma^{\prime}(1))\in[0,1]\times\{q\},\ t\in[0,1]$
    (we move along the bottom back curve in Fig. 6),
  3. (3)
    $(1,\gamma^{\prime}(1-t))\in\{1\}\times\mathbb{B}^{N},\ t\in[0,1]$
    (we move backward along the bottom red curve at the RHS in Fig. 6),
  4. (4)
    $(1-t,p=\gamma^{\prime}(0))\in[0,1]\times\{p\},\ t\in[0,1]$
    (we move backward along the upper back curve in Fig. 6).
We replace that path $F$ by a family $F[s_{1},s_{2}]$ of closed curves inside $[0,1]\times\mathbb{B}^{N}$ for any parameters $(s_{1},s_{2})\in[0,1]\times[0,1]$, namely we define $F[s_{1},s_{2}]$ by
  1. (1)
    $(0,\gamma^{\prime}(s_{1}t))\in\{0\}\times\mathbb{B}^{N},\ t\in[0,1]$,
    (we travel now only part of the way along the bottom red curve of the left hand side in Fig. 6)
  2. (2)
    $(s_{2}t,\gamma^{\prime}(s_{1}))\in[0,s_{2}]\times\{\gamma^{\prime}(s_{1})\},\ t\in[0,1]$,
    (we move parallel to the bottom back curve in Fig. 6, and in fact again only part of the way),
  3. (3)
    $(s_{2},\gamma^{\prime}(s_{1}(1-t)))\in\{s_{2}\}\times\mathbb{B}^{N},\ t\in[0,1]$,
    (we move now backwards parallel to the bottom red curve of the right hand side in Fig. 6),
  4. (4)
    $(s_{2}(1-t),p=\gamma^{\prime}(0))\in[0,s_{2}]\times\{p\},\ t\in[0,1]$
    (we move backwards along the upper back curve in Fig. 6).
The path $F=F(1,1)$ is null-homotopic by first deforming $F(1,1)$ to $F(0,1)$, and then deforming $F(0,1)$ to $F(0,0)\equiv\{(0,p)\}$. This proves (4.9). This proves that $\mathscr{J}^{\prime}$ commutes with the pullback connections.
We arrive at the following sequence of bundle isomorphisms (note that $r_{1}=\textup{id}$)
$\exp^{*}\mathscr{E}=r_{1}^{*}\exp^{*}\mathscr{E}\stackrel{{\scriptstyle\mathscr{J}^{\prime}}}{{\cong}}r_{0}^{*}\exp^{*}\mathscr{E}=(\exp\,\circ\,r_{0})^{*}\mathscr{E}=(\pi_{N}\circ r_{0})^{*}\mathscr{E}=\pi_{N}^{*}\mathscr{E}.$
@@
We can now define the Hilsum-Skandalis replacement for $f^{*}$. This requires a Thom form $\tau[{}^{e}TN]$ of the edge tangent bundle ${}^{e}TN$ as an additional ingredient. By definition, it is a closed differential form with compact support and of degree $\dim N$ on the total space of ${}^{e}TN$, such that its fibrewise integration gives the constant function $1$.
Hence, in a local trivialization, the Thom form is a family of top degree compactly supported differential forms along the fibres of ${}^{e}TN$, varying smoothly in the base $\widetilde{N}$. Up to rescaling, we may assume that $\tau[{}^{e}TN]$ is compactly supported in the disc bundle $\mathbb{B}^{N}$. Writing $p:{}^{e}TN\to\widetilde{N}$ for the bundle projection, we choose a connection that defines a decomposition into vertical and horizontal bundles and thus we get an isomorphism
${}^{e/w}T\mathbb{B}^{N}\cong p^{*}({}^{e/w}TN)\oplus T^{V}\mathbb{B}^{N},$
where the vertical component $T^{V}\mathbb{B}^{N}$ is the same both for the edge and wedge tangent bundles, since the fibres in $\mathbb{B}^{N}$ are open discs. As a differential form on $\mathbb{B}^{N}$, the Thom form $\tau[{}^{e}TN]$ has no horizontal components and hence can be viewed as a section of the exterior algebra of both ${}^{e}T^{*}\mathbb{B}^{N}$ and ${}^{w}T^{*}\mathbb{B}^{N}$, i.e.
$\tau[{}^{e}TN]\in C^{\infty}(\mathbb{B}^{N},\Lambda^{*}({}^{e/w}T^{*}\mathbb{B}^{N}).$
Thus, also the pullbacks of $\tau[{}^{e}TN]$ by (${}^{e}(\mathbb{B}\widetilde{f})^{*}$ and (${}^{w}(\mathbb{B}\widetilde{f})^{*}$ coincide, independent of whether we use edge or wedge pullback, i.e.
$(\mathbb{B}\widetilde{f})^{*}\tau[{}^{e}TN]:={}^{e}(\mathbb{B}\widetilde{f})^{*}\tau[{}^{e}TN]={}^{w}(\mathbb{B}\widetilde{f})^{*}\tau[{}^{e}TN]\in C^{\infty}(\widetilde{f}^{*}(\mathbb{B}^{N}),\Lambda^{*}({}^{e/w}T^{*}\widetilde{f}^{*}(\mathbb{B}^{N})).$
We can now define the Hilsum-Skandalis maps by a sequence of the following operations in each degree $k$
$C^{\infty}\Bigl{(}\widetilde{N},\Lambda^{k}({}^{e/w}T^{*}N)\otimes\mathscr{E}\Bigr{)}$
($\xrightarrow{{}^{e/w}(\exp\circ\,\mathbb{B}\widetilde{f})^{*}}\quad C^{\infty}\Bigl{(}\widetilde{f}^{*}(\mathbb{B}^{N}),\Lambda^{k}({}^{e/w}T^{*}\widetilde{f}^{*}(\mathbb{B}^{N}))\otimes(\exp\circ\,\mathbb{B}\widetilde{f})^{*}\mathscr{E}\Bigr{)}$
$\xrightarrow{\qquad\mathscr{J}\qquad}\quad C^{\infty}\Bigl{(}\widetilde{f}^{*}(\mathbb{B}^{N}),\Lambda^{k}({}^{e/w}T^{*}\widetilde{f}^{*}(\mathbb{B}^{N}))\otimes(\widetilde{f}\circ\pi)^{*}\mathscr{E}\Bigr{)}$
$\xrightarrow{(\mathbb{B}\widetilde{f})^{*}\tau\bigl{[}{}^{e}TN\bigr{]}\wedge}\quad C^{\infty}\Bigl{(}\widetilde{f}^{*}(\mathbb{B}^{N}),\Lambda^{k+\dim N}({}^{e/w}T^{*}\widetilde{f}^{*}(\mathbb{B}^{N}))\otimes(\widetilde{f}\circ\pi)^{*}\mathscr{E}\Bigr{)}$
$\xrightarrow{\qquad\pi_{*}\qquad}\quad C^{\infty}\Bigl{(}\widetilde{M},\Lambda^{k}({}^{e/w}T^{*}M)\otimes\widetilde{f}^{*}\mathscr{E}\Bigr{)}.$